OpenStudy (anonymous):

A 25 kg rock is on the edge of a 125 m cliff. What potential energy does the rock possess? If the rock falls from the cliff, how fast is it moving just before it strikes the ground?

OpenStudy (anonymous):

P.E=mgh. =25*125*9.8 =30625 K.E=P.E (o.5)mv^2=mgh v=(30625*2)^(.5) v=247.4873m/s